# A Data-Centric Perspective on Evaluating Machine Learning Models for Tabular Data

### v2

@@ 4.3 Feature Engineering is Still the Most Important Factor for Top Performance
The most remarkable performance gains are achieved through feature engineering.    Figure 6 shows that expert feature engineering is the most important modeling component on average. This holds true for all models, indicating that unlike for modalities like imaging, neural networks do not automate feature engineering for tabular data. When comparing the performance of different models in the standardized preprocessing pipeline (blue/orange/green bars in Figure 3), we can observe that using any other model than CatBoost rarely brings large gains. Only for the SCS dataset does FTTransformer clearly outperform all other models. For all other datasets, the average performance gains achievable solely with model selection are small. Hence, our results confirm the findings of McElfresh et al., [63] that model selection is less important than HPO on a strong tree-based baseline for most datasets. Furthermore, we extend this finding by quantifying the even more important aspect of dataset-specific feature engineering.
@@
The optimal treatment of categorical features can be dataset-specific.    Table 2 shows that a different treatment of categorical features than the model-inherent treatment was necessary for two datasets to achieve top performance. Furthermore, each of the two datasets required a different encoding method. This shows that standardized preprocessing can be biased for categorical features. Hence, whenever the goal is not to evaluate models as AutoML solutions, categorical data treatment methods in comparative studies should not only be model-specific, but also dataset-specific.
@@ 4.4 The Importance of Test-Time Adaptation and Temporal Characteristics
Models in real-world applications are often applied to non-i.i.d. tabular data.    By definition, TTA should only be effective if the data violates the i.i.d. assumption and contains distribution shifts to adapt to. Indeed, the data collection process likely happened over time for most of the datasets used in our framework. However, timestamps were not always provided as the competitions were conceptualized as static tabular data tasks. Therefore, most of the datasets were also used in at least one comparative study for tabular data, although non-i.i.d. was a criterion for exclusion (e.g., SVPC, AEAC, and PSSDP in [30], SCTP and OGPCC in [47], or MBGM in [35]). Our results show, that despite treating datasets as static, the samples remain non-i.i.d. and approaches like test-time adaptation can improve performance. Furthermore, there is evidence that other datasets treated as i.i.d. in related work actually have a temporal nature. I.e., Kohli et al., [47] found that the most frequently used dataset in tabular data research is the forest cover type dataset [12]. At the same time, this dataset is used as a benchmark in online learning to measure the ability of models to adapt to concept drifts [24, 73]. As models for tabular data assume the data to be i.i.d., most benchmarks for evaluating tabular general-purpose models either directly name the data being non-i.i.d. as an exclusion criterion [35, 29] or exclude data that requires special CV procedures [11], which leads to the same results. In contrast, our analysis of Kaggle competitions revealed that most tabular data competitions have temporal characteristics and that the best solutions for such datasets typically engineer time-invariant features and utilize tabular data models assuming the data to be i.i.d (i.e. [3]). We conclude that there might be a mismatch between current evaluation frameworks for tabular data in academia and the tabular data tasks practitioners were interested in getting solved through ML competitions on Kaggle.
@@ 5 Implications for Future Work
Dinstinguish between AutoML and model comparisons.    Our findings highlight that standardized evaluation setups do not necessarily ensure fair model comparisons. In standardized preprocessing setups, models are treated as AutoML solutions, whereas in real-world applications, they are components of highly dataset-specific pipelines. Comparative studies should differentiate between datasets that benefit from known feature engineering steps and those that are self-sufficient. One approach could be to separate raw data benchmarks from fully preprocessed benchmarks, as done in our study. Standardized setups are more suitable for benchmarking AutoML solutions, while feature-engineered setups might be better for benchmarking models. Future research could emphasize incorporating dataset-specific (expert) preprocessing pipelines into benchmarks. However, gathering high-quality expert solutions at a large scale is tedious and may require a community effort.
@@
Improve feature engineering capabilities of models for tabular data. Researchers developing general-purpose models should recognize the impact of feature engineering on model performance. CatBoost has advanced the field by automating feature engineering on categorical data. However, significant feature engineering effort is still necessary for datasets where this is not the only challenge. Future work should take a data-centric perspective and focus on automating successful feature engineering techniques through novel architecture components. Our expert feature engineering pipelines can serve as a starting point for evaluating and developing new methods. Furthermore, unlike previously claimed [35], categorical features can indeed be an important challenge for deep learning models. Hence, future work could focus on improvements over classic embeddings for categorical features in general-purpose deep learning architectures.
Methods for tabular data with temporal characteristics. Our analysis highlights the importance of distribution shifts in real-world tabular data, even when treating a task as static. Future work could investigate test-time adaptation methods specifically for tabular data, using our datasets and the identified test-time feature engineering techniques as baselines. Furthermore, our findings indicate that the current research focus on static i.i.d. data might hinder the development of techniques to handle weak temporal correlations in tabular data. Future work should focus on developing models with inductive biases for tabular data with temporal characteristics.
@@ Appendix A Datasets and Expert Solutions
In this Section, we provide more details on the dataset/competition selection process and the expert solutions implemented in our framework. Table 4 shows the name of all Kaggle competitions included in our framework as well as a hyperlink for the implemented expert feature engineering.
@@ A.1 Dataset Selection
Name Competition Name Implemented Expert Solution
MBGM mercedes-benz-greener-manufacturing 1st place solution
SVPC santander-value-prediction-challenge 6th place solution
AEAC amazon-employee-access-challenge 1st place solution
OGPCC otto-group-product-classification-challenge 8th place solution
SCS santander-customer-satisfaction 3rd place solution
BPCCM bnp-paribas-cardif-claims-management 8th place solution
SCTP santander-customer-transaction-prediction 1st place solution
HQC homesite-quote-conversion 15th place solution
IFD ieee-fraud-detection 1st place solution
PSSDP porto-seguro-safe-driver-prediction 2nd place solution
Table 4: Datasets and expert solutions included in our framework. The competitions can be accessed at https://www.kaggle.com/competitions/{Competition Name}.
@@ A.2 Implemented Components of Expert Solutions
  • Occurrence of numeric values from train data in the test data (SCTP)
@@
Cross-validation procedures.    We used the same CV split type for each dataset as the expert solutions but unified the number of folds across all our datasets. We used 10 folds for most datasets, as this worked well for all datasets. We used 10-fold cross-validation for all but one dataset. For classification tasks, the folds were stratified across the target. For the IFD dataset, we split the data based on the month the data was collected, which resulted in six folds. For faster training, fewer folds could be used for large datasets with similar results. For large datasets, most expert solutions used fewer folds, e.g., for the PSSDP competition.
@@ A.3 Discussion on Test-Time Feature Engineering
We argue that the common ML setup allowing for test-time adaptation corresponds to a frequent real-world application scenario where 1) The data to predict arrives in batches, 2) No real-time predictions are required, and 3) Retraining the model at test time is feasible. The first criterion is important as the employed test-time feature engineering techniques (e.g. dimensionality reduction and frequency encoding) often required the presence of many test samples at once. It is unclear whether this kind of domain adaptation would work per sample. The second and third criteria are necessary as test-time feature engineering always requires retraining the model, which is infeasible in online applications. In contrast, test-time feature engineering is not applicable in applications where online predictions are required, the number of test samples is small, or not retrainable (e.g., large-scale) models are used. Importantly, while test-time feature engineering might be infeasible in such applications, other test-time adaptation techniques might still apply. One example of such a task conceptualization amenable to test-time feature engineering is product return prediction, where samples are collected over a day, and a (lightweight) model can be retrained daily. In this scenario, using the test data in an unsupervised fashion for better adaptation to possible distribution shifts is feasible. After examining the application scenarios of the tasks in our framework, we found that most of them, like customer transaction prediction (SCTP) and customer satisfaction prediction (SCS), would allow such a setup, although likely with smaller amounts of test data than used for the competitions. Furthermore, our discussion in Subsection 4.4 reveals that many tabular datasets not in our scope have temporal components and thus may be amenable to test-time feature engineering.
@@ Appendix C Detailed Performance Results
In this Section, we provide the leaderboard position results for all the experiments in the main paper, separated by hyperparameter regime and preprocessing pipeline. The results can be seen in Tables 13, 14, and 16.
@@ Appendix D Additional Results
D.1 Comparison of Preprocessing Pipelines per Model and Dataset

### v3

@@ 4.3 Feature Engineering is Still the Most Important Factor for Top Performance
The most remarkable performance gains are achieved through feature engineering.    Figure 6 shows that expert feature engineering is the most important modeling component on average. This holds true for all models, indicating that unlike for modalities like imaging, neural networks do not automate feature engineering for tabular data. When comparing the performance of different models in the standardized preprocessing pipeline (blue/orange/green bars in Figure 3), we can observe that using any other model than CatBoost rarely brings large gains. Only for the SCS dataset does FTTransformer clearly outperform all other models. For all other datasets, the average performance gains achievable solely with model selection are small.
@@
Hence, our results confirm the findings of McElfresh et al., [55] that model selection is less important than HPO on a strong tree-based baseline for most datasets. Furthermore, we extend this finding by quantifying the even more important aspect of dataset-specific feature engineering.
@@
The optimal treatment of categorical features can be dataset-specific.    Table 2 shows that a different treatment of categorical features than the model-inherent treatment was necessary for two datasets to achieve top performance. Furthermore, each of the two datasets required a different encoding method. This shows that standardized preprocessing can be biased for categorical features. Furthermore, the performance of deep learning models on categorical data can be greatly improved with feature engineering techniques on categorical data. I.e., for the AEAC dataset, which consists entirely of categorical features, all neural networks gain from feature engineering techniques like categorical feature interactions, as can be seen in Figure 3. This suggests that current architectures do not adequately capture the complex patterns within categorical data. Hence, whenever the goal is not to evaluate models as AutoML solutions, categorical data treatment methods in comparative studies should not only be model-specific, but also dataset-specific.
@@ 4.4 The Importance of Test-Time Adaptation and Temporal Characteristics
Models in real-world applications are often applied to non-i.i.d. tabular data.    By definition, TTA should only be effective if the data violates the i.i.d. assumption and contains distribution shifts to adapt to. Indeed, the data collection process likely happened over time for most of the datasets used in our framework. However, timestamps were not always provided as the competitions were conceptualized as static tabular data tasks. Therefore, most of the datasets were also used in at least one comparative study for tabular data, although non-i.i.d. was a criterion for exclusion (e.g., SVPC, AEAC, and PSSDP in [21], SCTP and OGPCC in [41], or MBGM in [26]). Our results show, that despite treating datasets as static, the samples remain non-i.i.d. and approaches like test-time adaptation can improve performance. Furthermore, there is evidence that other datasets treated as i.i.d. in related work actually have a temporal nature. I.e., the electricity dataset [29] is frequently used in academic benchmarks [26, 55]. At the same time, this dataset would actually require a time-based data split and is used as a benchmark in online learning to measure the ability of models to adapt to concept drifts [15]. As models for tabular data assume the data to be i.i.d., most benchmarks for evaluating tabular general-purpose models either directly name the data being non-i.i.d. as an exclusion criterion [26, 21] or exclude data that requires special CV procedures [4], which leads to the same results. In contrast, our analysis of Kaggle competitions revealed that most tabular data competitions have temporal characteristics and that the best solutions for such datasets typically engineer time-invariant features and utilize tabular data models assuming the data to be i.i.d (i.e. [32]). We conclude that there might be a mismatch between current evaluation frameworks for tabular data in academia and the tabular data tasks practitioners were interested in getting solved through ML competitions on Kaggle.
4.5 Limitations
Despite the outlined advantages our evaluation has some limitations compared to evaluation designs in academic benchmarks:
  1. 1.
    The distribution of leaderboard rankings varies across competitions, and advancements in one competition’s leaderboard do not necessarily translate equivalently in another. Moreover, large gains on a leaderboard may correspond to only minor increases on the actual task metric, particularly in competitions where solutions are widely shared and numerous similar entries are submitted. As an alternative, we repeat our evaluation based on the original task metric in Appendix F.
  2. 2.
    Not all leaderboard submissions are made by experts, resulting in a tail of lower-quality entries in each competition. Users of our framework should exercise caution when interpreting the leaderboard as a performance metric, ensuring that only factually accurate statements about its implications are used.
  3. 3.
    Due to the extent of our experiments, it was infeasible to repeat them to obtain error bars. This limitation is important, as it leaves the randomness introduced by data splits, model-specific characteristics, and hyperparameter optimization unquantified. Hence, small differences between models on single datasets need to be interpreted with caution. Nevertheless, our primary focus was to assess the effects of various modeling components. In this regard, the extensive nature of our experiments and the clear distinctions observed between preprocessing pipelines across multiple models and datasets suggest a minimal impact of randomness on our main findings. For users applying our framework to systematically compare individual models, alternative configurations with repeated evaluations would be necessary to increase reliability.
  4. 4.
    Due to the focus on incentivized competitions, most datasets are from the finance domain and primarily represent North America and Europe, leading to an underrepresentation of other domains and regions. To address this limitation, future analyses could incorporate competitions from additional platforms. However, as our framework was designed to be user-friendly, we focused on Kaggle, which contains an API for effortlessly downloading datasets and submitting predictions.
A more extensive discussion of limitations can be found in Appendix E.
@@ 5 Implications for Future Work
More careful choice of preprocessing for model evaluation.    Our findings highlight that standardized evaluation setups do not necessarily ensure fair model comparisons. In standardized preprocessing setups, models are evaluated as if they were AutoML solutions, whereas in real-world applications, they are components of highly dataset-specific pipelines. Researchers should be aware a) whether their datasets are amenable to feature engineering, b) that standardized preprocessing setups treat models as AutoML systems, and, c) that true ceteris paribus (c.p.) comparisons are hard if some models (i.e., CatBoost) apply feature engineering internally and others don’t. Feature-engineered evaluation setups can be more suitable if a study aims for truly c.p. conditions or for evaluating models in realistic scenarios. Standardized evaluation setups are more suitable for benchmarking AutoML solutions and can also be suitable if models are expected to learn features without human effort. Future research could emphasize incorporating dataset-specific (expert) preprocessing pipelines into benchmarks or separate raw data benchmarks from fully preprocessed benchmarks, as done in our study. However, gathering high-quality expert solutions at a large scale is tedious and may require a community effort.
@@
Investigate why some feature engineering operations are not inherently learned by models. Researchers developing general-purpose models should recognize the impact of feature engineering on model performance. CatBoost has advanced the field by automating feature engineering on categorical data. However, significant feature engineering effort is still necessary for datasets where this is not the only challenge. Our study made it evident that there are transformations of the feature space which are not learned by models without manual feature engineering. While we focused on a pipeline perspective, future work could look at particular feature engineering techniques to uncover modes of failure for current models and develop novel architecture components. I.e., our experiments show that deep learning models benefit from feature engineering on categorical features. Hence, unlike previously claimed [26], categorical features can indeed be an important challenge for deep learning models and future work could focus on improvements over conventional embeddings. Our expert feature engineering pipelines can serve as a starting point for evaluating and developing new methods. Furthermore, AutoGluon was sometimes outperformed by single models in our analysis, although it contains the same models. Future work could investigate why AutoGluon does not always benefit from feature engineering to the same extent as single models.
Methods for tabular data with temporal characteristics. Our analysis highlights the temporal nature of many real-world tabular data tasks, as well as the importance of accounting for distribution shifts. Future work could investigate test-time adaptation methods specifically for tabular data, using our datasets and the identified test-time feature engineering techniques as baselines. Furthermore, our findings indicate that the current research focus on static i.i.d. data might hinder the development of techniques to handle weak temporal correlations in tabular data. Future work should focus on developing models with inductive biases for tabular data with temporal characteristics.
@@
Acknowledgments and Disclosure of Funding
This research was supported by the Ministry of Economic Affairs, Labour and Tourism Baden-Württemberg. Furthermore, the authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. We want to thank the reviewers of the NeurIPS datasets and benchmarks track for their invaluable feedback, which greatly improved the quality of this paper. Moreover, we want to thank Kaggle for hosting the competitions used in our framework and maintaining them as open source. In addition, we thank the Kaggle users gmobaz, joeytaj, pyduan, bensolucky, mikeskim, leustagos, efimov, titericz, davutpolat, jiweiliu, kazanova, stasg7, mmueller, kyazuki, utility, raddar, confirm, psilogram, fl2ooo, konradb, rejulien, cdeotte, kyakovlev, xiaozhouwang, kelexu, and qqgeogor for sharing the competition solutions we used. Lastly, we thank Malte Gaber and Michael Temnov for their assistance in implementing Kaggle solutions.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Each main claim is supported in one of the Subsections of Section 4
    2. (b)
      Did you describe the limitations of your work? [Yes] Limitations are mentioned in the main paper where appropriate and discussed in more detail in the Appendix.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] There are no direct societal impacts.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] The paper does not include theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] The paper does not include theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] The training details are provided in the supplementary material.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Due to the extent of our experiments (over 200,000 trained models), repeating experiments to obtain error bars was not feasible.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] The used hardware is described in the supplementary material.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We greatly rely on datasets and public solutions from Kaggle competitions and reference them in Table 1 and the supplementary material.
    2. (b)
      Did you mention the license of the assets? [Yes] All assets are publicly available.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Our code is provided to the reviewers in the supplementary material and will be made publicly available upon acceptance.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] All utilized data is publicly available.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] All utilized data was anonymized by the publishers.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Datasets and Expert Solutions
In this Section, we provide more details on the dataset/competition selection process and the expert solutions implemented in our framework. Table 4 shows the name of all Kaggle competitions along with additional information.
@@ A.1 Dataset Selection
Name Competition Name End date Expert solution openml-id
MBGM mercedes-benz-greener-manufacturing 2017-06-11 1st place 42570
SVPC santander-value-prediction-challenge 2018-08-21 6th place 42572
AEAC amazon-employee-access-challenge 2013-08-01 1st place 4135
OGPCC otto-group-product-classification-challenge 2015-05-19 8th place 45548
SCS santander-customer-satisfaction 2016-05-03 3rd place -
BPCCM bnp-paribas-cardif-claims-management 2016-04-19 8th place -
SCTP santander-customer-transaction-prediction 2019-04-11 1st place 42395
HQC homesite-quote-conversion 2016-02-09 15th place -
IFD ieee-fraud-detection 2019-10-04 1st place -
PSSDP porto-seguro-safe-driver-prediction 2017-11-30 2nd place 43121
Table 4: Datasets and expert solutions included in our framework. The competitions can be accessed at https://www.kaggle.com/competitions/{Competition Name}. The openml id is provided for better contextualization with prior work.
@@ A.2 Implemented Components of Expert Solutions
  • Occurrence of numeric features from train data in the test data (SCTP)
@@
Cross-validation procedures.    We used the same CV split type for each dataset as the expert solutions but unified the number of folds across all our datasets. We used 10 folds for most datasets, as this worked well for all datasets. For classification tasks, the folds were stratified across the target. For the IFD dataset, we split the data based on the month the data was collected, which resulted in six folds. For faster training, fewer folds could be used for large datasets with similar results. For large datasets, most expert solutions used fewer folds, e.g., for the PSSDP competition.
@@ A.3 Discussion on Test-Time Feature Engineering
We argue that the common ML setup allowing for test-time adaptation corresponds to a frequent real-world application scenario where 1) The data to predict arrives in batches, 2) No real-time predictions are required, 3) The train data is still available at test time, and 4) Retraining the model at test time is feasible. The first criterion is important as the employed test-time feature engineering techniques (e.g. dimensionality reduction and frequency encoding) often required the presence of many test samples at once. It is unclear whether this kind of domain adaptation would work per sample. The other criteria are necessary as test-time feature engineering always requires retraining the model. Consequently, test-time feature engineering is not applicable in online learning, if the number of test samples is small, or if the models are not retrainable (i.e., large-scale models). Importantly, while test-time feature engineering might be infeasible in such applications, other test-time adaptation techniques might still apply. One example of a task conceptualization amenable to test-time feature engineering is product return prediction, where samples are collected over a day, and a (lightweight) model can be retrained daily. In this scenario, using the test data in an unsupervised fashion for better adaptation to possible distribution shifts is feasible. After examining the application scenarios of the tasks in our framework, we found that most of them, like customer transaction prediction (SCTP) and customer satisfaction prediction (SCS), would allow such a setup, although likely with smaller amounts of test data than used for the competitions. Furthermore, our discussion in Subsection 4.4 reveals that many tabular datasets not in our scope have temporal components and thus may be amenable to test-time feature engineering.
@@ Appendix C Detailed Performance Results
In this Section, we provide the leaderboard position results for all the experiments in the main paper, separated by hyperparameter regime and preprocessing pipeline. The results can be seen in Tables 13, 14, 15, and 16.
@@ Appendix D Additional Results
D.1 Runtime analysis
As our experiments were conducted with varying hardware setups depending on availability, we cannot directly compare runtimes. Hence, we conduct a separate experiment where we train each model with default hyperparameters on each dataset and preprocessing pipeline on the same hardware. In particular, we train all models in a setup where all folds are trained in parallel distributed equally over two NVIDIA RTX A6000 GPUs. This setup is particularly beneficial for small datasets as well as for MLP-like neural architectures as sequential training does not fully utilize the capabilities of modern GPUs on these datasets and models. Figure 7 shows that XGBoost is the only Pareto point after feature engineering and with test-time adaptation as it is the fastest and best performing model at the same time. For the standardized setup the tree-based models build the Pareto frontier. However, as in this plot runtime is measured with default hyperparameters and performance with extensively tuned hyperparameters, the results do not represent the actual time required to obtain the performance results.
Figure 7: Runtime analysis. Performance is reported as the leaderboard position after extensive hyperparameter optimization averaged over the ten datasets. Time is reported as the total time in minutes required to train a model with default hyperparameters on all ten datasets with a particular preprocessing pipeline. Note that each experiment was conducted with parallelizing the training of folds, s.t. the time in GPU hours is much higher than the reported time. The black line represents the Pareto frontier for the models trained in the standardized preprocessing pipeline. For the other pipelines, XGBoost was the only Pareto optimum.
D.2 Model Ranking Variances Across Dataset
Figure 17 shows that the dataset-variance is rather high, mainly because the hardness of the task differs among datasets, especially with standardized preprocessing. It can be seen that the variance of the best performing methods strongly decreases after TTA, as good preprocessing eases the prediction tasks for the previously hard datasets.
CatBoost XGBoost LightGBM MLP-PLR FTTransformer GRANDE ResNet
Stand. 0.79 (0.21) 0.68 (0.17) 0.62 (0.19) 0.6 (0.22) 0.59 (0.19) 0.52 (0.23) 0.44 (0.21)
FE 0.89 (0.14) 0.9 (0.15) 0.85 (0.19) 0.82 (0.19) 0.81 (0.2) 0.81 (0.17) 0.62 (0.27)
TTA 0.97 (0.04) 0.97 (0.03) 0.9 (0.16) 0.91 (0.12) 0.91 (0.1) 0.83 (0.16) 0.64 (0.28)
Table 17: Average leaderboard position with variance per dataset and pipeline.
D.3 Comparison of Preprocessing Pipelines per Model and Dataset